OpenStudy (anonymous):

what is x+3y=-6 in slope intercept form?

OpenStudy (anonymous):

y=-x/3-2

OpenStudy (anonymous):

you have to get y by itself so you subtract x and you get 3y=-x-6 then you divide the 3 leaving you with y=-x/3-2

OpenStudy (anonymous):

thanks!